# ShowTable: Unlocking Creative Table Visualization with Collaborative Reflection and Refinement

### v1

@@ A.2 Reinforcement Learning for Image Generation
Diffusion models have established themselves as the predominant framework for text-to-image (T2I) generation [flux, sd, podell2023sdxl, ddpm, qwenimage]. The integration of reinforcement learning (RL) into this paradigm began with works utilizing policy gradient optimization to guide the denoising process [ddpo, rewarddance, imagereward, dpok]. The field subsequently expanded to include preference-based alignment methods, which achieve competitive performance without explicit reward modeling [dpo]. A significant recent development is the adoption of Group Relative Policy Optimization(GRPO) [deepseekr1, deepseekmath], an efficient alternative that has inspired numerous adaptations for T2I generation. These include pioneering works [dancegrpo, flowgrpo] which unified diffusion and flow matching under an SDE-based formulation. This line of inquiry further explores the design of specialized reward models and data curation strategies to enhance the framework’s capability for producing high-quality, preference-aligned visual outputs [mixgrpo, prefgrpo, branchgrpo].
@@ B.2 Rewriting Training Details
As discussed in Section 3.2, we fine-tune a specialized rewriting module based on Qwen3-8B to handle the critical task of reasoning and compositional planning. This module is trained on our 30K SFT data with both thinking and rewriting result, and we construct various kinds of instruction templates during training to ensure the diversity. For the implementation, we utilize the LLaMA-Factory [llamafactory] library. We train the model for 3 epochs with a total batch size of 256. The training employs a learning rate of 1e-5, combined with a cosine learning rate decay strategy to stabilize the training process.
@@ B.3 Refinement Training Details
To empower the refinement model with the capability for precise, fine-grained edits on dense data infographics, we employ an on-policy reinforcement learning approach utilizing the Group Relative Policy Optimization (GRPO) [deepseekr1] algorithm. In this framework, rendering accuracy serves as the critical reward signal.
@@
To address this, we develop a specialized, efficient Reward Model (RM). We construct a pairwise dataset $D$ consisting of positive and negative graphic samples generated from the same prompt, as detailed in Section 4.2. We fine-tune a Qwen2.5-VL-3B [qwen2.5vl] model, denoted as $f_{\theta}$, to serve as the quality assessor. For a given prompt $p$, with $x_{w}$ denoting the preferred (positive) sample and $x_{l}$ the dispreferred (negative) sample, the model is optimized using the Bradley-Terry (BT) loss [bradley1952bradterry]:
@@
where $\sigma(\cdot)$ represents the Sigmoid function. To stabilize the output, the reward score is computed by extracting and averaging the probabilities corresponding to the tokens for digits 0–9 from the output logits. The final reward signal used in RL is a weighted combination: $R=0.8\cdot f_{\theta}(x,p)+0.2\cdot\text{ImageReward}(x,p)$ [imagereward].
@@
Implementation. Following the Flow-GRPO framework [flowgrpo], we train our refinement model using a distilled 8-step version of Qwen-Image-Edit-2509 [qwenimagelightning] to accelerate training efficiency. The model is trained on our constructed 5K refinement dataset for 1 epoch using 32 GPUs. We set the image resolution to $1024\times 1024$ and perform 16 rollouts per prompt. We utilize 8 sampling steps for both the inference rollout and the training backward pass. As illustrated in Figure A3, compared to the unstable baseline using raw LLM scores, our approach with the trained Reward Model achieves stable and consistent performance gains.
@@ Appendix D More Experiments
In the main paper, we primarily focused on open-source models to ensure reproducibility. Here, we extend our evaluation to the recently released Wan2.5-Preview series [wan2.5] to assess the ShowTable pipeline’s performance with state-of-the-art (SOTA) generation capabilities. The results are presented in Table A1.

### v2

@@ A.2 Reinforcement Learning for Image Generation
Diffusion models have established themselves as the predominant framework for text-to-image (T2I) generation [5, 46, 44, 23, 65, 29] and text-to-video (T2V) generation [30, 61, 64, 62]. The integration of reinforcement learning (RL) into this paradigm began with works utilizing policy gradient optimization to guide the denoising process [4, 67, 71, 16]. The field subsequently expanded to include preference-based alignment methods, which achieve competitive performance without explicit reward modeling [57]. A significant recent development is the adoption of Group Relative Policy Optimization(GRPO) [21, 47], an efficient alternative that has inspired numerous adaptations for T2I generation. These include pioneering works [72, 37] which unified diffusion and flow matching under an SDE-based formulation. This line of inquiry further explores the design of specialized reward models and data curation strategies to enhance the framework’s capability for producing high-quality, preference-aligned visual outputs [32, 59, 33].
@@ B.2 Rewriting Training Details
As discussed in Section 3.2, we fine-tune a specialized rewriting module based on Qwen3-8B to handle the critical task of reasoning and compositional planning. This module is trained on our 30K SFT data with both thinking and rewriting result, and we construct various kinds of instruction templates during training to ensure the diversity. For the implementation, we utilize the LLaMA-Factory [79] library. We train the model for 3 epochs with a total batch size of 256. The training employs a learning rate of 1e-5, combined with a cosine learning rate decay strategy to stabilize the training process.
@@ B.3 Refinement Training Details
To empower the refinement model with the capability for precise, fine-grained edits on dense data infographics, we employ an on-policy reinforcement learning approach utilizing the Group Relative Policy Optimization (GRPO) [21] algorithm. In this framework, rendering accuracy serves as the critical reward signal.
@@
To address this, we develop a specialized, efficient Reward Model (RM). We construct a pairwise dataset $D$ consisting of positive and negative graphic samples generated from the same prompt, as detailed in Section 4.2. We fine-tune a Qwen2.5-VL-3B [2] model, denoted as $f_{\theta}$, to serve as the quality assessor. For a given prompt $p$, with $x_{w}$ denoting the preferred (positive) sample and $x_{l}$ the dispreferred (negative) sample, the model is optimized using the Bradley-Terry (BT) loss [6]:
@@
where $\sigma(\cdot)$ represents the Sigmoid function. To stabilize the output, the reward score is computed by extracting and averaging the probabilities corresponding to the tokens for digits 0–9 from the output logits. The final reward signal used in RL is a weighted combination: $R=0.8\cdot f_{\theta}(x,p)+0.2\cdot\text{ImageReward}(x,p)$ [71].
@@
Implementation. Following the Flow-GRPO framework [37], we train our refinement model using a distilled 8-step version of Qwen-Image-Edit-2509 [42] to accelerate training efficiency. The model is trained on our constructed 5K refinement dataset for 1 epoch using 32 GPUs. We set the image resolution to $1024\times 1024$ and perform 16 rollouts per prompt. We utilize 8 sampling steps for both the inference rollout and the training backward pass. As illustrated in Figure A4, compared to the unstable baseline using raw LLM scores, our approach with the trained Reward Model achieves stable and consistent performance gains.
@@ Appendix D More Experiments
In the main paper, we primarily focused on open-source models to ensure reproducibility. Here, we extend our evaluation to the recently released Wan2.5-Preview series [51] to assess the ShowTable pipeline’s performance with state-of-the-art (SOTA) generation capabilities. The results are presented in Table A2.
@@ D.2 More Visualization Results
Table A3: User study results showing human preference rates (%) across five evaluation dimensions. Our full pipeline achieves the highest human preference in almost all dimensions, particularly dominating in Data Accuracy (DA).
Method DA TR $R$ AA AQ
Qwen-Image 2.25 38.50 12.00 2.50 7.75
RW + Qwen-Image 24.00 29.25 41.25 48.25 44.50
RW + Qwen-Image + REF 73.75 32.25 46.75 49.25 47.75
D.3 User Study
To further validate the effectiveness of our proposed pipeline from a human perspective, we conducted a comprehensive user study. We randomly selected 20 table instances and generated visualizations using three configurations: the base generation model (Qwen-Image), the model prefixed with the rewriting module (RW+Qwen-Image), and our full ShowTable pipeline (RW + Qwen-Image + REF). We invited 20 human evaluators to independently assess the 20 sets of images. Participants were asked to vote for the best generation in each set across our five evaluation dimensions. The preference rates are reported in LABEL:tab:user_study. Consistent with our quantitative automated metrics, the full ShowTable pipeline significantly outperforms the baselines in human preference. Notably, ShowTable achieves an overwhelming 73.75% preference rate in Data Accuracy, confirming that our progressive reflection-refinement loop effectively corrects data-to-visual mapping errors that humans easily perceive. For Text Rendering (TR), the base model occasionally receives higher preference (38.50%); this aligns with our observation, where the base model often defaults to merely rendering the raw table text directly without executing complex layout reasoning or data visualization. While this “lazy rendering” avoids text artifacts, it completely fails the core visualization task (reflected in its dismal 2.25% DA score). Overall, the human study strongly corroborates that our pipeline produces the most factually faithful and visually appealing infographics.